OpenStudy (anonymous):

3r – 5t – 4rt r=-4 t=3

OpenStudy (anonymous):

i think you are supposed to replace \(r\) by \(-4\) and \(t\) by \(3\) which will give you \[3\times (-4)-5\times 3-4\times (-4)\times 3\]

OpenStudy (anonymous):

then some arithmetic to get the actual number

OpenStudy (anonymous):

idk how to do it though